OpenStudy (anonymous):

A ship sails 20 miles east of port then 15 miles south to an island. How far is the ship from the port if you measured in a straight line? Choose from one of these answers: 5,10,20,25

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

use pythogoras theorem for rt angled trianlgle and hence the ans=sqrt(225+400)=25

OpenStudy (anonymous):

\[\sqrt{25^{}}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you